OpenStudy (anonymous):

Find the value of X if 180=(12x-22)2

OpenStudy (anonymous):

180=24x-44 24x=180+44 24x =224 x=224\24

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

thanks! i really appreciate it!!!! :) enjoy ur medal!